# CoCoA: A Generalized Approach to Uncertainty Quantification by Integrating Confidence and Consistency of LLM Outputs

### v2

@@ 3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
We start by summarizing the benefits and drawbacks of various uncertainty quantification approaches discussed above:
  1. 1.
    Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures.
  2. 2.
    Output probabilities $p(\mathbf{y}^{(j)}\mid\mathbf{x}),\,j=1,\dots,M$ might have substantially different values for semantically equivalent outputs, which questions the usefulness of (weighted) averaging these probabilities for uncertainty quantification.
  3. 3.
    For various methods based on the aggregation over multiple samples, the result might be suboptimal due to the noise related to the averaging over all generated outputs. Focusing solely on a particular output sequence and its relation to other generated outputs might be beneficial.
@@
In what follows, we present a family of UQ Confidence and Consistency-based Approaches (CoCoA), offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs.
Let us consider an actual output sequence $\mathbf{y}_{*}$ and a set of sampled sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$. Here, $\mathbf{y}_{*}$ might be one of the sequences $\mathbf{y}^{(i)}$ or might be generated separately. In what follows, we will consider several possible cases, including $\mathbf{y}_{*}$ being a random sequence from a set $\{\mathbf{y}^{(i)}\}$, $\mathbf{y}_{*}$ being a sequence from a set $\{\mathbf{y}^{(i)}\}$ having the highest probability, and, finally, $\mathbf{y}_{*}$ being a sequence found via the beam search procedure.
First, consider an information-based uncertainty score of the output $\mathbf{y}_{*}$:
$u_{*}^{\text{info}}=U^{\text{info}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (7)
where $U^{\text{info}}$ might be MSP, perplexity, mean token entropy, or another uncertainty measure related solely to the generated sequence $\mathbf{y}_{*}$.
We quantify the consistency-based uncertainty via a direct measurement of the semantic similarity of generated sequence $\mathbf{y}_{*}$ to sampled sequences:
$u_{*}^{\text{cons}}=\frac{1}{M}\sum_{i=1}^{M}(1-g_{*i}),$ (8)
where $g_{*i}=g\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$. This formulation satisfies the desired properties of the uncertainty function – that is their values are nonnegative and their values increase with increased inconsistency (decreasing value of $g_{*i}$). In our ablation study, we will show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix B.3).
Finally, we need to aggregate $u_{*}^{\text{info}}$ and $u_{*}^{\text{cons}}$ into a single uncertainty measure. We propose to aggregate them in a multiplicative way:
@@
$u_{*}^{\texttt{CoCoA}}=u_{*}^{\text{info}}\cdot u_{*}^{\text{cons}}$ (9)
@@
This formulation preserves the non-negativity and the monotonicity properties while integrating both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences that are both intrinsically uncertain (high $u_{*}^{\text{info}}$) and semantically inconsistent with the dataset (high $u_{*}^{\text{cons}}$), while keeping it low for the opposite scenario (see Figure 2).
Although the choice of the multiplicative aggregation function CoCoA is heuristic, it provides a practical and effective way to combine information- and consistency-based uncertainty signals in LLMs. In our ablation study, we also compare the multiplicative formulation in (9) to a simpler additive variant, $u_{*}^{\text{info}}+u_{*}^{\text{cons}}$ (see Appendix B.3). Empirically, the multiplicative combination is better at capturing the joint impact of both information-based and consistency-based uncertainty, yielding more reliable estimation across all tasks.
@@ 4.1 Experimental Setup
To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2024; Fadeeva et al., 2023) by implementing our approach within its framework. Since the library already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), summarization (SUM), and machine translation (MT), compared to existing baselines.
@@
Datasets.
For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (Joshi et al., 2017), an open-domain factual QA dataset; CoQA (Reddy et al., 2019), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (Hendrycks et al., 2021), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (Cobbe et al., 2021), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (Bojar et al., 2014) and WMT19 German-English (Barrault et al., 2019). Finally, for summarization, we used XSum (Narayan et al., 2018), a dataset of complex documents paired with concise abstractive summaries.
For all datasets, we follow (Vashurin et al., 2024) for selecting the subsets, for prompt formatting, and for number and sourcing of few-shot examples.
Models.
@@
We evaluated our method using the base versions of three open-weights language models: LLaMA 3.1 8B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Falcon 3 7B (Team, 2024). The open-source nature of these models provides full access to their token probabilities, which are essential for implementing our UQ method. For all models, we consider base versions, without instruction-tuning.
Metric Llama Mistral Falcon
@@
MCSE 0.357 0.380 0.192 0.453 0.406 0.162 0.460 0.409 0.128

### v3

@@ 3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures. There exist approaches that bridge the gap between information-based and consistency-based methods that show great promise but lack the fundamental base; see discussion in Section 4. In what follows, we present a family of Confidence and Consistency-based Approaches (CoCoA) for UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs via the unified MBR-based framework.
If we consider only the consistency measure, like the one in (8), we omit the information that is contained in the model confidence. Thus, we want the resulting uncertainty measure to explicitly consider both semantic consistency and model-based uncertainty. Let us consider the risk of the form:
$r(\mathbf{y},\mathbf{y}^{\prime}\mid\mathbf{x})=u(\mathbf{y}\mid\mathbf{x})\cdot\bigl{(}1-s(\mathbf{y},\mathbf{y}^{\prime})\bigr{)},$ (9)
where
  • $s(\mathbf{y},\mathbf{y}^{\prime})\in[0,1]$ is any utility function used in the standard MBR decoding (usually representing semantic similarity);
  • $u(\mathbf{y}\mid\mathbf{x})\geq 0$ is the model-based uncertainty measure for the output sequence $\mathbf{y}$. For example, we can use $u(\mathbf{y}\mid\mathbf{x})=1-p(\mathbf{y}\mid\mathbf{x})$ (similarly to (6)) or $u(\mathbf{y}\mid\mathbf{x})=-\log p(\mathbf{y}\mid\mathbf{x})$.
Then, the corresponding Bayes risk is $R_{\text{CoCoA}}(\mathbf{y}\mid\mathbf{x})=u(\mathbf{y}\mid\mathbf{x})\cdot\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\bigl{(}1-s(\mathbf{y},\mathbf{y}^{\prime})\bigr{)}$ and for an output sequence $\mathbf{y}_{*}$, the resulting uncertainty measure becomes:
$U_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\bigl{(}1-s(\mathbf{y}_{*},\mathbf{y}^{\prime})\bigr{)}.$ (10)
Finally, given a set of samples $\bigl{\{}\mathbf{y}^{(i)}\bigr{\}}_{i=1}^{M}$ we obtain an empirical uncertainty estimate:
@@
$\widehat{U}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (11)
where $s_{*i}=s\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$.
@@
The resulting uncertainty measure integrates both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences $\mathbf{y}_{*}$ that are both intrinsically uncertain (high $u(\mathbf{y}_{*}\mid\mathbf{x})$) and semantically inconsistent with the other samples (high $\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i})$), while keeping it low for the opposite scenario; see Figure 2.
CoCoA Light. Computing the consistency-based uncertainty measure $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ within $\widehat{U}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})$ requires multiple samples, leading to significant computational overhead. To address this, we propose to approximate the behavior of $\widehat{U}_{\text{cons}}(\mathbf{x})$ with a learned function. Our method closely matches the original sampling-based measure in quality while requiring only greedy output generation during inference, eliminating the need for additional sampling. Notably, this approximation can be learned without access to ground-truth labels, relying solely on the input data and the associated uncertainty values.
Thus, the CoCoA uncertainty measure that incorporates an approximation of the consistency uncertainty $\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})\approx\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ can be defined as:
$\widehat{U}^{L}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (12)
We name this approach CoCoA Light. Below, we describe how this approximation can be obtained.
Assume access to a held-out set of input sequences $\mathbf{x}_{j},j=1,\dots,n$. We emphasize that this held-out set is not labeled. For each sequence, we extract the embeddings $\{e(\mathbf{y}_{j}\mid\mathbf{x}_{j})\}_{j=1}^{n}$ of the corresponding greedy model generations $\mathbf{y}_{j}$. The corresponding targets are the consistency uncertainty scores ${\widehat{U}_{\text{cons}}(\mathbf{y}_{j}\mid\mathbf{x}_{j})},j=1,\dots,n$, computed via Monte Carlo estimation of the minimum Bayes risk; see equation (8).
A lightweight auxiliary model, $\widehat{g}\bigl{(}e(\mathbf{y}_{*}\mid\mathbf{x})\bigr{)}$, is trained in a supervised fashion to map model embeddings to uncertainty scores. During inference, the model generates a greedy output $\mathbf{y}_{\text{test}}$ for a test input $\mathbf{x}_{\text{test}}$, from which the embeddings $e(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})$ are obtained and passed to the auxiliary model to predict the uncertainty score:
$\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})=\widehat{g}\bigl{(}e(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})\bigr{)}.$ (13)
We note that a similar approach (Kossen et al., 2024) has been previously proposed to approximate semantic entropy using hidden states from the model.
4 Related Work
Information-based methods. These methods are one of the most commonly used. They quantify uncertainty by analyzing the probability distributions of tokens within a given output. The simplest of these methods, Maximum Sequence Probability (MSP), measures the probability of the most likely output sequence given a specific input. Perplexity is another common uncertainty measure: it gauges how well the model predicts each token in a sequence and is formally defined as the exponential of the mean negative log-likelihood of those tokens (Fomicheva et al., 2020). The Mean Token Entropy method evaluates token-level predictions across the entire sequence by computing the average entropy of the token probability distributions at each position (Fomicheva et al., 2020). While these methods provide useful uncertainty estimates, they do not address the broader uncertainty inherent in generative tasks, as they rely on a single sequence and fail to capture the diversity of possible outputs that the model could generate for the same input.
Consistency-Based Methods. These methods estimate uncertainty by generating multiple sequences from the model’s output distribution and analyzing the variability among the sampled outputs. These methods are particularly relevant to NLP tasks, as they account for scenarios in which multiple plausible outputs may exist for a given input. Among the simplest consistency-based methods are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian Lin et al. (2024), which quantify uncertainty by how many distinct “meanings” the model produces. While effective at capturing global variation, they do not yield uncertainty scores for individual responses. To overcome this, the diagonal of the Degree Matrix was proposed to assess how similar each output is to the rest, enabling per-response uncertainty estimation Lin et al. (2024). A related approach, Lexical Similarity, operates by calculating the average similarity of words or phrases across every pair of responses in the sample Fomicheva et al. (2020).
Information-Based Methods with Repeated Sampling. More recent methods attempt to reconcile output consistency with information-based signals—i.e., they not only look at how varied outputs are, but also how probable each output is under the model’s own generation process. For instance, Semantic Entropy (Kuhn et al., 2023) groups outputs into semantically homogeneous clusters—capturing distinct “meanings” that might just differ in phrasing—and calculates entropy across these clusters. SentenceSAR refines this idea by defining “relevance” as the sum of pairwise similarities between a given sentence and others weighted by each sentence’s model-assigned generation probability (Duan et al., 2024). SAR combines the sentence-level relevance of SentenceSAR with token-level probability adjustments, attempting a more fine-grained balance of semantic and token-level uncertainty (Duan et al., 2024). Semantic Density, Qiu and Miikkulainen (2024) evaluates uncertainty by evaluating how densely a model’s generated response is situated within the semantic space of all possible outputs, with lower density indicating higher uncertainty. One limitation of sampling-based approaches is their computational cost: generating multiple outputs from the model can be expensive. Recent work has sought to address this by learning to predict semantic uncertainty directly from hidden states, allowing for the approximation of semantic entropy without repeated sampling Kossen et al. (2024). Finally, BSDetector (Chen and Mueller, 2024) proposes an additive composition of self-reported confidence with observed consistency, however it relies on questionable ability of the model to assess it’s own confidence in text form. Another limitation is that it requires selection of a trade-off coefficient between confidence and consistency parts.
While these probability-weighted measures can provide deeper insights, they also sometimes underperform in practice (Vashurin et al., 2025), especially when the model’s probability estimates are unreliable, or when important nuances of semantic diversity are lost in the weighting process. Proper balancing raw output consistency with generation likelihood remains an open problem.
Minimum Bayes Risk for LLMs. Recent work has proposed incorporating uncertainty into Minimum Bayes Risk decoding by introducing a posterior over model parameters, enabling uncertainty-aware generation Daheim et al. (2025). Complementing this, a Bayesian framework for subjective uncertainty quantification was developed yielding improved calibration in generation tasks (Wang and Holmes, 2025).
@@ 5.1 Experimental Setup
To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2025; Fadeeva et al., 2023) by implementing our approach within its framework. Since the library already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), summarization (SUM), and machine translation (MT), compared to existing baselines.
Datasets. For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (Joshi et al., 2017), an open-domain factual QA dataset; CoQA (Reddy et al., 2019), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (Hendrycks et al., 2021), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (Cobbe et al., 2021), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (Bojar et al., 2014) and WMT19 German-English (Barrault et al., 2019). Finally, for summarization, we used XSUM (Narayan et al., 2018), a dataset of complex documents paired with concise abstractive summaries.
For all datasets, we follow (Vashurin et al., 2025) for selecting the subsets, for prompt formatting, and for the number and sourcing of few-shot examples.
@@
Models. We evaluated our method using the base versions of three open-weight language models: LLaMA 3.1 8B Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and Falcon 3 7B Team (2024). The open-source nature of these models provides full access to their token probabilities, which are essential for implementing our UQ method. For all models, we consider base versions, without instruction tuning.
Similarity Function. To measure the similarity, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (Liu et al., 2019; Reimers and Gurevych, 2019; Cer et al., 2017). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure.
Appendix C.1 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
Baselines. We compare the performance of the proposed method against a diverse set of baselines and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. Information-based methods include Maximum Sequence Probability (MSP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). Consistency-based methods include the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, hybrid methods include Semantic Entropy and SAR. All formulations for these baselines can be found in Appendix D. Lastly, we evaluate performance of $\hat{U}_{\text{cons}}$ (Consistency) and $\hat{U}_{\text{cons}}^{L}$ (Consitency Light) as an uncertainty measure.
Evaluation measure. As our evaluation measure, we choose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (Malinin and Gales, 2021). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold $a$ and observing how the average quality $Q(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of the remaining predictions changes (Figure 3). The metric is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates quality measures.
For some of the datasets we also report AUROC results, see Appendix G.
Quality Measures. PRR requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between the correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content and the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
To further improve comprehensiveness of reported results, we considered alternative choices of quality measures, and report corresponding PRR scores in Appendix F.
Generation Setup. We discuss the generation parameters, the decoding strategy, and sample selection procedure in depth in Appendix A. In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with a focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
CoCoA Light Details. The multilayer perceptron is used as an auxiliary model. The features used are embeddings from the middle layers of the base LLM, which are the most informative according to recent work Chen et al. (2024). In particular, for Llama 3.1 8B and Mistral 7B models, embeddings are taken from the 16th layer, and for Falcon 3 7B from the 14th layer.
Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
@@
MCSE 0.310 0.323 0.033 0.389 0.304 0.007 0.414 0.317 0.159
MCNSE 0.309 0.393 0.022 0.384 0.410 0.009 0.405 0.422 0.108
DegMat 0.406 0.302 0.081 0.423 0.305 0.137 0.483 0.353 0.201
SAR 0.414 0.455 0.077 0.462 0.435 0.094 0.481 0.458 0.144
Consistency 0.437 0.421 0.024 0.471 0.392 0.051 0.494 0.416 0.226
Consistency Light 0.390 0.458 -0.022 0.444 0.387 -0.006 0.427 0.476 0.232
MSP 0.409 0.399 0.328 0.475 0.383 0.287 0.475 0.356 0.201
$\text{CoCoA}_{MSP}\,\,\text{Light}$ 0.449 $\uparrow$ 0.502 $\uparrow$ 0.358 $\uparrow$ 0.503 $\uparrow$ 0.480 $\uparrow$ 0.309 $\uparrow$ 0.492 $\uparrow$ 0.514 $\uparrow$ 0.242 $\uparrow$
PPL 0.381 0.386 0.369 0.424 0.427 0.204 0.456 0.450 0.155
$\text{CoCoA}_{PPL}\,\,\text{Light}$ 0.445 $\uparrow$ 0.487 $\uparrow$ 0.382 $\uparrow$ 0.479 $\uparrow$ 0.480 $\uparrow$ 0.260 $\uparrow$ 0.512 $\uparrow$ 0.528 $\uparrow$ 0.234 $\uparrow$
MTE 0.353 0.382 0.357 0.417 0.438 0.182 0.456 0.473 0.152
$\text{CoCoA}_{MTE}$ 0.447 $\uparrow$ 0.478 $\uparrow$ 0.380 $\uparrow$ 0.492 $\uparrow$ 0.469 $\uparrow$ 0.288 $\uparrow$ 0.527 $\uparrow$ 0.508 $\uparrow$ 0.228 $\uparrow$
$\text{CoCoA}_{MTE}\,\,\text{Light}$ 0.428 $\uparrow$ 0.494 $\uparrow$ 0.372 $\uparrow$ 0.475 $\uparrow$ 0.482 $\uparrow$ 0.254 $\uparrow$ 0.507 $\uparrow$ 0.533 $\uparrow$ 0.234 $\uparrow$
5.2 Results
Main results. Table 1 shows the PRR scores under the greedy generation setup (See Appendix B for Best Sample and MBR Sample). We report aggregated PRR for each type of task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix E.
We can see that our CoCoA methods are the best across all tasks and models. They outperform existing consistency-based and hybrid state-of-the-art approaches, like Semantic Entropy and SAR. In addition, the proposed CoCoA approach consistently surpasses the baseline UQ measures: for example, $\texttt{CoCoA}_{PPL}$ outperforms standard Perplexity, illustrating the advantage of combining token-level confidence with semantic consistency. This pattern holds for other information-based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
Ablation study. As a part of our ablation study, we investigate the impact of different similarity measures (see Appendix C.1). We find that for some tasks, the similarity score computed by the Cross-encoder does not yield optimal performance. For example, for question-answering tasks on CoQA and TriviaQA, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
The next section of our ablation study focuses on alternative forms of combining model confidence $u(\mathbf{y}_{*}\mid\mathbf{x})$ and consistency $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$; see Appendix C.2. First, we consider an additive form of combining them:
$\widehat{U}_{\text{AdditiveCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})+\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}).$ (15)
The results show that this additive formulation does not perform as well as the multiplicative one. The additive form tends to underemphasize the interaction between the two components, which is critical for capturing the nuanced relationships between confidence and consistency.
We also consider an alternative formulation of the consistency term $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$, as the average of the full pairwise dissimilarity. In this formulation, $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples. Our experiments demonstrate that this formulation is not very strong. By distributing the consistency computation across all samples, it loses focus on the specific sequence being evaluated.
Lastly, in Appendix C.2, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty estimation. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
6 Limitations
Task and Domain Dependency. CoCoA relies on an information-based confidence score and a semantic similarity function, whose effectiveness can vary across models, tasks, and domains. In open-ended tasks like creative generation, producing diverse outputs is expected; on the other hand, tasks requiring precise reasoning can be sensitive to subtle errors that generic similarity metrics may miss. Adapting these components to specific domains remains an important direction for future work.
Limited Sample Size. CoCoA estimates consistency by sampling multiple outputs. Generating many samples is computationally expensive, so our experiments—like most sampling-based methods—use relatively small sets. While even a few samples can yield meaningful estimates, they may miss the full diversity of model outputs for complex prompts.
Quality Metric. Finally, the CoCoA’s performance assessment depends on quality metrics (e.g., COMET for machine translation, and Accuracy for QA) that may not capture every nuance of textual outputs. Further refining or extending quality metrics to account for deeper reasoning, factual faithfulness, and stylistic appropriateness would better align uncertainty scores with real-world perceptions of model correctness.
7 Conclusion
We presented CoCoA, a unified approach that integrates Confidence and Consistency for uncertainty quantification in LLMs. By combining confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone. In extensive evaluations on question answering, summarization, and translation, our approach outperformed existing baselines and state-of-the-art UQ methods. Moreover, CoCoA’s flexible design allows easy adaptation to a variety of tasks and settings.
Appendix A Decoding Strategy and Sample Selection
To facilitate the choice of decoding and sample selection strategies for our experiments, we conducted an evaluation of model performance with different approaches to both. Table 2 shows average values of corresponding quality metrics for all combinations of models and datasets. We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation:
  • Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus no further selection of sample is needed.
  • Most Probable sample selects the output with highest model-assigned probability among several sampled outputs.
  • MBR sample selects the output with highest average consistency with respect to other sampled outputs.
We note that selecting a random sample from the model outputs incurs a significant drop in the quality of results on several datasets, most prominently on GSM8k. Based on these observations, we evaluate the efficacy of UQ in three setups: greedy decoding, stochastic sampling with a focus on the most probable sample and MBR decoding.
Regardless of the way the main model response was obtained, responses that were used to quantify consistency in (8) were generated via repeated stochastic sampling. In all the experiments, where stochastic sampling is involved, it was performed with temperature $T=1.0$, top-k equal to 50, and top-p equal to $1.0$.
Sample Metric Greedy Random Sample Most Probable Sample MBR Sample
Falcon7b-Base
CoQA Align Score 0.793 0.706 0.783 0.793
GSM8k Accuracy 0.776 0.313 0.205 0.419
MMLU Accuracy 0.715 0.638 0.715 0.699
TriviaQA Align Score 0.557 0.473 0.568 0.559
WMT14FrEn Comet 0.867 0.833 0.857 0.855
WMT19DeEn Comet 0.846 0.807 0.826 0.828
XSUM Align Score 0.842 0.734 0.782 0.801
Llama8b-Base
CoQA Align Score 0.756 0.671 0.743 0.766
GSM8k Accuracy 0.548 0.234 0.261 0.346
MMLU Accuracy 0.570 0.368 0.577 0.469
TriviaQA Align Score 0.686 0.625 0.687 0.699
WMT14FrEn Comet 0.863 0.819 0.852 0.844
WMT19DeEn Comet 0.870 0.816 0.854 0.845
XSUM Align Score 0.848 0.608 0.825 0.703
Mistral7b-Base
CoQA Align Score 0.793 0.695 0.777 0.790
GSM8k Accuracy 0.382 0.169 0.190 0.259
MMLU Accuracy 0.632 0.552 0.632 0.612
TriviaQA Align Score 0.743 0.655 0.750 0.739
WMT14FrEn Comet 0.863 0.812 0.830 0.841
WMT19DeEn Comet 0.864 0.805 0.836 0.837
XSUM Align Score 0.803 0.578 0.775 0.665
Appendix B Results Summary for Most Probable Sample and MBR Sample strategies
Tables 3 and 4 report the PRR scores under the Most Probable Sample and MBR Sample generation setups, as discussed in Section  5.2. We observe that CoCoA-family methods, and their Supervised versions, consistently improve base uncertainty estimators.
Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
QA NMT SUM QA NMT SUM QA NMT SUM
MCSE 0.356 0.380 0.192 0.453 0.406 0.162 0.460 0.409 0.128